# A complete classification of the zero-dimensional homogeneous spaces under determinacy

### v2

@@ 4.2. Relativizable pointclasses and universal sets
Let $S\in\mathbf{\Pi}(2^{\omega}\times Z)$ be a $2^{\omega}$-universal set for $\mathbf{\Pi}(Z)$. Using [Ke, Theorem 13.6], we can fix an embedding $j:2^{\omega}\longrightarrow Z$. Set $W=j[2^{\omega}]$. Notice that $(j\times\mathsf{id}_{Z})[U]\in\mathbf{\Pi}(W\times Z)$ because $\mathbf{\Pi}$ is a topological pointclass. Since $\mathbf{\Pi}$ is relativizable, there exists $T\in\mathbf{\Pi}(Z\times Z)$ such that $T\cap(W\times Z)=(j\times\mathsf{id}_{Z})[S]$. It is straightforward to check that $T$ is a $Z$-universal set for $\mathbf{\Pi}(Z)$. ∎
Lemma 4.13.
Let $\mathbf{\Pi}$ be a topological pointclass, and let $Z$ be a space. If there exists a $Z$-universal set for $\mathbf{\Pi}(Z)$ then $\mathbf{\Pi}(Z)$ is non-selfdual.
Proof.
Fix a $Z$-universal set $S$ for $\mathbf{\Pi}(Z)$. Assume, in order to get a contradiction, that $\mathbf{\Pi}(Z)$ is selfdual. Define $f:Z\longrightarrow Z\times Z$ by setting $f(z)=(z,z)$ for $z\in Z$, and observe that $f$ is continuous. It follows that $f^{-1}[S]\in\mathbf{\Pi}(Z)$, hence $Z\setminus f^{-1}[S]\in\mathbf{\Pi}(Z)$. Therefore, since $S$ is $Z$-universal, we can fix $z\in Z$ such that $S_{z}=Z\setminus f^{-1}[S]$. It is straightforward to check that $z\in S_{z}$ iff $z\notin S_{z}$, which is a contradiction. ∎
Finally, we are in a position to characterize the induced relativizable pointclasses, under suitable determinacy assumptions. Notice that Theorem 4.15 can be viewed as an improved version of Theorem 4.7.
@@
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $\mathbf{\Pi}$ be a relativizable pointclass such that $\mathbf{\Pi}(Z)\subseteq\mathbf{\Sigma}(Z)$ for every zero-dimensional Borel space $Z$. Then the following conditions are equivalent:
  1. (1)
    $\mathbf{\Pi}$ is induced by some $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$,
  2. (2)
    There exists $D\subseteq\mathcal{P}(\omega)$ such that $\mathbf{\Pi}(Z)=\mathbf{\Gamma}_{D}(Z)$ for every zero-dimensional Borel space $Z$,
  3. (3)
    There exists $D\subseteq\mathcal{P}(\omega)$ such that $\mathbf{\Pi}(Z)=\mathbf{\Gamma}_{D}(Z)$ for some uncountable zero-dimensional Borel space $Z$,
  4. (4)
    $\mathbf{\Pi}(Z)$ has a $2^{\omega}$-universal set for every zero-dimensional Borel space $Z$,
  5. (5)
    $\mathbf{\Pi}(Z)$ has a $2^{\omega}$-universal set for some uncountable zero-dimensional Polish space $Z$,
  6. (6)
    $\mathbf{\Pi}(Z)\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$ set for every uncountable zero-dimensional Polish space $Z$,
  7. (7)
    $\mathbf{\Pi}(Z)\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$ for some zero-dimensional Polish space $Z$.
Proof.
The implication $(1)\rightarrow(3)$ follows from Theorem 4.7, and the implication $(2)\rightarrow(3)$ is trivial. In order to prove that $(3)\rightarrow(2)$, assume that $(3)$ holds for some uncountable zero-dimensional Borel space $Z$. Pick a zero-dimensional Borel space $W$, and assume without loss of generality that $W\subseteq Z$. Then
$\mathbf{\Pi}(W)=\{A\cap W:A\in\mathbf{\Pi}(Z)\}=\{A\cap W:A\in\mathbf{\Gamma}_{D}(Z)\}=\mathbf{\Gamma}_{D}(W),$
where the first equality holds because $\mathbf{\Pi}$ is relativizable, and the last equality follows from Lemma 4.6.
The implication $(2)\rightarrow(4)$ follows from Proposition 4.11, and the implication $(4)\rightarrow(5)$ is trivial. To see that the implication $(5)\rightarrow(6)$ holds, fix an uncountable zero-dimensional Polish space $Z$ such that $\mathbf{\Pi}(Z)$ has a $2^{\omega}$-universal set. Notice that $\mathbf{\Pi}(Z)$ is continuously closed because $\mathbf{\Pi}$ is a topological pointclass. Furthermore, Lemmas 4.12 and 4.13 guarantee that $\mathbf{\Pi}(Z)$ is non-selfdual. It follows from Lemma 3.5 that $\mathbf{\Pi}(Z)=\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$, hence $\mathbf{\Pi}(Z)=\mathbf{\Gamma}(Z)$ for some $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$. Now let $W$ be an uncountable zero-dimensional Polish space. Since $Z$ is uncountable, we can assume without loss of generality that $W\subseteq Z$. Then
$\mathbf{\Pi}(W)=\{A\cap W:A\in\mathbf{\Pi}(Z)\}=\{A\cap W:A\in\mathbf{\Gamma}(Z)\}=\mathbf{\Gamma}(W),$
where the first equality holds because $\mathbf{\Pi}$ is relativizable, and the last equality follows from Lemma 3.12. Since $W$ is uncountable, it follows from Theorem 3.13 that $\mathbf{\Pi}(W)\in\mathsf{NSD}_{\mathbf{\Sigma}}(W)$, as desired.
The implication $(6)\rightarrow(7)$ is trivial. Therefore, in order to conclude the proof, it will be enough to show that $(7)\rightarrow(1)$. So fix an uncountable zero-dimensional Polish space $Z$ such that $\mathbf{\Pi}(Z)\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$.
Claim. $\mathbf{\Pi}(\omega^{\omega})\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$.
Proof of Claim. By [Ke, Theorem 7.8], we can assume without loss of generality that $Z$ is a closed subspace of $\omega^{\omega}$. By [Ke, Proposition 2.8], we can fix a retraction $\rho:\omega^{\omega}\longrightarrow Z$. By Lemma 3.5, it will be enough to show that $\mathbf{\Pi}(\omega^{\omega})$ is non-selfdual. Assume, in order to get a contradiction, that $\mathbf{\Pi}(\omega^{\omega})$ is selfdual, and pick $A$ such that $A\!\!\downarrow\,=\mathbf{\Pi}(Z)$. Then $\rho^{-1}[Z\setminus A]=\omega^{\omega}\setminus\rho^{-1}[A]\in\mathbf{\Pi}(\omega^{\omega})$, hence
$Z\setminus A=Z\cap(\rho^{-1}[Z\setminus A])\in\mathbf{\Pi}(Z)$
because $\mathbf{\Pi}$ is relativizable. This contradicts the fact that $\mathbf{\Pi}(Z)$ is non-selfdual. $\blacksquare$
Set $\mathbf{\Gamma}=\mathbf{\Pi}(\omega^{\omega})$. First, using arguments similar to those in the proof of the above claim, one can easily show that $\mathbf{\Pi}(W)=\mathbf{\Gamma}(W)$ for every zero-dimensional Polish space $W$. Then, using the fact that $\mathbf{\Pi}$ is relativizable and Lemma 3.12, it is straightforward to check that $\mathbf{\Pi}$ is induced by $\mathbf{\Gamma}$. ∎
4.3. New Wadge classes from old
Many examples of Wadge classes are obtained by combining elements of known classes using a fixed set-theoretic operation (consider for example the difference hierarchy). In fact, combining sets that belong to a given sequence of non-selfdual Wadge classes according to a given Hausdorff operation always yields a non-selfdual Wadge class (this follows from Corollary 4.16 by ignoring the sets $U_{2k+1}$). However, the operation of separated unions (to be introduced in Definition 6.5) will require in addition that certain sets are pairwise disjoint. This is why, in this section, the statements of our results look more complicated than one might expect.
We begin by showing that combining relativizable pointclasses according to a fixed Hausdorff operation preserves the property of having a universal set. Notice that Theorem 4.15 does not require any determinacy assumptions. The desired result on combining Wadge class will then follow easily, thanks to Theorem 4.14.
Theorem 4.15.
@@ 4.3. New Wadge classes from old
Let $D\subseteq\mathcal{P}(\omega)$, let $\mathbf{\Pi}_{n}$ for $n\in\omega$ be relativizable topological pointclasses, and let $\xi<\omega_{1}$. Assume that each $\mathbf{\Pi}_{n}$ has a $2^{\omega}$-universal set. Set
(4.1) $\mathbf{\Pi}(Z)=\{\mathcal{H}_{D}(A_{0},U_{1},A_{2},U_{3}\ldots):A_{2k}\in\mathbf{\Pi}_{k}(Z)\text{ for every }k\in\omega\text{,}\\ U_{2k+1}\in\mathbf{\Sigma}^{0}_{1+\xi}(Z)\text{ for every }k\in\omega\text{, and }U_{2i+1}\cap U_{2j+1}\text{ whenever }i\neq j\}$
@@
Using the assumption that each $\mathbf{\Pi}_{n}$ is a relativizable pointclass and Theorem 2.4, it is straightforward to check that $\mathbf{\Pi}$ is a relativizable pointclass. Pick a zero-dimensional space $Z$. Fix a homeomorphism $h:2^{\omega}\longrightarrow(2^{\omega})^{\omega}$. Set $(y)_{n}=\pi_{n}(h(y))$ for $y\in 2^{\omega}$ and $n\in\omega$, where $\pi_{n}:(2^{\omega})^{\omega}\longrightarrow 2^{\omega}$ denotes the projection on the $n$-th coordinate. By assumption, for every $k\in\omega$ we can fix a $2^{\omega}$-universal set $S_{2k}$ for $\mathbf{\Pi}_{k}(Z)$. Define
@@
By [Ke, Theorem 22.3], we can fix a $2^{\omega}$-universal set $U$ for $\mathbf{\Sigma}^{0}_{1+\xi}(Z)$. Define
@@
for $k\in\omega$, and observe that each $U_{2k+1}\in\mathbf{\Sigma}^{0}_{1+\xi}(2^{\omega}\times Z)$. By Theorem 2.4, there exist $U^{\prime}_{2k+1}\in\mathbf{\Sigma}^{0}_{1+\xi}(2^{\omega}\times Z)$ for $k\in\omega$ such that the following conditions are satisfied:
@@
will yield precisely the sequences consisting of pairwise disjoint elements of $\mathbf{\Sigma}^{0}_{1+\xi}(Z)$. Furthermore, changing the values of $(y)_{n}$ for even $n$ will not affect the above sequence.
@@
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $D\subseteq\mathcal{P}(\omega)$, let $Z$ be an uncountable zero-dimensional Polish space, and let $\xi<\omega_{1}$. If $\mathbf{\Gamma}_{n}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$ for $n\in\omega$ and
(4.2) $\mathbf{\Gamma}=\{\mathcal{H}_{D}(A_{0},U_{1},A_{2},U_{3}\ldots):A_{2k}\in\mathbf{\Gamma}_{n}\text{ for every }k\in\omega\text{,}\\ U_{2k+1}\in\mathbf{\Sigma}^{0}_{1+\xi}(Z)\text{ for every }k\in\omega\text{, and }U_{2i+1}\cap U_{2j+1}\text{ whenever }i\neq j\}$
@@
Pick $\mathbf{\Gamma}_{n}(Z)\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$ for $n\in\omega$, where each $\mathbf{\Gamma}_{n}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$. Let $\mathbf{\Pi}_{n}$ be the relativizable pointclass induced by $\mathbf{\Gamma}_{n}$ for $n\in\omega$. It is clear that $\mathbf{\Gamma}=\mathbf{\Pi}(Z)$, where $\mathbf{\Pi}$ is defined as in the statement of Theorem 4.15. Therefore, the desired conclusion follows from Theorems 4.14 and 4.15. ∎
@@ 4.4. Clarifying level and expansions
In this section, we will show how to view level and expansions through the lens of Hausdorff operations. While the only result from this section that will be needed in the rest of the article is Proposition 4.19, the alternative descriptions given here are helpful in clarifying these two important concepts.
@@
As a first application, we will give some concrete examples of expansions. This will be useful in Section 7.2.
@@
We conclude this section with a characterization of the level that is perhaps more intuitive than the official Definition 3.23. The Borel version of Theorem 4.20 originally appeared as part of [LSR1, Théorème 8].
@@
Chapter 5 Wadge theory: the separation property
The purpose of this chapter is to introduce the separation property, and to give a complete proof that exactly one Wadge class in every non-selfdual pair has the separation property (see Theorem 5.8). This theorem is essentially due to Steel and Van Wesep, and it will be important for our existence results, because of the role that it plays in the proof of Theorem 7.18, which in turn will be needed to obtain one of the closure properties of Section 7.5. It will also be applied directly in the proofs of Corollaries 9.15 and 9.17. However, the separation property is not relevant to the uniqueness theorem.
5.1. Definition and auxiliary results
Much like the socks in a pair, a non-selfdual Wadge class $\mathbf{\Gamma}$ and its dual $\widecheck{\mathbf{\Gamma}}$ do not appear to be distinguishable in any obvious way. Luckily, work of Steel and Van Wesep provides an elegant solution to this problem, at least when the ambient space is $\omega^{\omega}$. The key notion is the (first) separation property, already mentioned in the introduction and formally defined below.111 The second separation property will only be needed in this section and the next, and it will not appear anywhere else in this article. According to [VW1, page 92] and [VW2, page 77], both separation properties are due to Lusin. More precisely, they showed that exactly one of $\mathbf{\Gamma}$ and $\widecheck{\mathbf{\Gamma}}$ has the separation property. We will show that their result holds for uncountable zero-dimensional Polish spaces (see Theorem 5.8).
@@ 5.1. Definition and auxiliary results
We conclude this section by describing a method that is at the heart of the results of the next two sections.222 According to [VW1, page 99] and [VW2, page 81], this method was invented by Martin (see the proof of [VW3, Theorem 2.5]).
Lemma 5.3 (Martin).
Let $\Omega$ be a non-empty discrete space, and let $\sigma_{i}$ be a strategy for Player I on $\omega$ for $i\in\Omega$. Set $f_{i}=f_{\sigma_{i}}$ for $i\in\Omega$. Then there exists a function $\lambda:\omega\times\Omega^{\omega}\longrightarrow\omega^{\omega}$ satisfying the following conditions, where we use the notation $z^{\ast}_{n}=\lambda(n,z)$:
  1. (1)
    $f_{z(n)}(z^{\ast}_{n+1})=z^{\ast}_{n}$ for every $z\in\Omega^{\omega}$ and $n\in\omega$,
  2. (2)
    $(i^{\frown}z)^{\ast}_{n+1}=z^{\ast}_{n}$ for every $z\in\Omega^{\omega}$, $i\in\Omega$ and $n\in\omega$,
  3. (3)
    $(i^{\frown}z)^{\ast}_{0}=f_{i}(z^{\ast}_{0})$ for every $z\in\Omega^{\omega}$ and $i\in\Omega$,
  4. (4)
    $\lambda$ is continuous.
Proof.
First, we give a slightly informal description of the construction of $\lambda$, based on the following diagram. Assuming that $z\in\Omega^{\omega}$ has been fixed, the $n$-th element of the first row (starting from the left) is simply the first element of $\omega$ that Player I should play according to $\sigma_{z(n)}$. Then, one obtains the second row by letting $z^{\ast}_{n}(1)$ be the response of Player I to $z^{\ast}_{n+1}(0)$ according to $\sigma_{z(n)}$. By repeating this process, one obtains all the remaining rows.
[…]
$\ddots$
[…]
More formally, the definition of $\lambda$ is given by requiring that
$z^{\ast}_{n}\upharpoonright(k+1)=\sigma_{z(n)}(z^{\ast}_{n+1}\upharpoonright k)$
for every $z\in\Omega^{\omega}$ and $n,k\in\omega$. Using this equality, one sees that
$f_{z(n)}(z^{\ast}_{n+1})=\bigcup_{k\in\omega}\sigma_{z(n)}(z^{\ast}_{n+1}\upharpoonright k)=\bigcup_{k\in\omega}z^{\ast}_{n}\upharpoonright(k+1)=z^{\ast}_{n}$
for every $z\in\Omega^{\omega}$ and $n\in\omega$. Therefore $(1)$ holds. Furthermore, given $z\in\Omega^{\omega}$, $i\in\Omega$ and $n\in\omega$, it is straightforward to prove by induction that $(i^{\frown}z)^{\ast}_{n+1}\upharpoonright k=z^{\ast}_{n}\upharpoonright k$ for each $k$. Hence $(2)$ holds. To prove that $(3)$ holds, simply observe that
$(i^{\frown}z)^{\ast}_{0}=f_{i}((i^{\frown}z)^{\ast}_{1})=f_{i}(z^{\ast}_{0})$
for every $z\in\Omega^{\omega}$ and $i\in\Omega$, where the first equality holds by $(1)$ and the second equality holds by $(2)$. Given $(n,z)\in\omega\times\Omega^{\omega}$ and $k\in\omega$, one sees that $\lambda(n,z^{\prime})\upharpoonright k=\lambda(n,z)\upharpoonright k$ whenever $z^{\prime}\in\omega^{\omega}$ and $z^{\prime}(i)=z(i)$ for all $i\in\omega$ such that $n\leq i<n+k$. This shows that $(4)$ holds. ∎
5.2. At most one class in every pair has the separation property
All the results in this section are essentially due to Van Wesep, except that we have given more general formulations, with the goal of obtaining Theorem 5.8. Lemmas 5.4 and 5.5 are “extracted” from the proof of [VW1, Theorem 4.2.1] (see also [VW2, Theorem 1] and [VW3, Theorem 5.2]), while Theorem 5.6 is taken almost verbatim from [VW1, Theorem 4.2.2] (see also [VW2, Theorem 2] and [VW3, Theorem 5.3]), except that we have given a more systematic account of Martin’s method (see Lemma 5.3).
@@ 5.2. At most one class in every pair has the separation property
By [Ke, Corollary 6.5], we can fix a subspace $K$ of $Z$ such that $K\approx 2^{\omega}$. Let $h:2^{\omega}\times 2^{\omega}\longrightarrow K$ be a homeomorphism. Define $j:2^{\omega}\times 2^{\omega}\times Z\longrightarrow Z\times Z$ by setting $j(x,y,z)=(h(x,y),z)$ for $(x,y,z)\in 2^{\omega}\times 2^{\omega}\times Z$, and observe that $j$ is an embedding with $\mathsf{ran}(j)=K\times Z$. By Lemma 3.12, there exist $A_{0},A_{1}\in\mathbf{\Gamma}(Z\times Z)$ such that $A_{0}\cap(K\times Z)=j[A^{\prime}_{0}]$ and $A_{1}\cap(K\times Z)=j[A^{\prime}_{1}]$. It is straightforward to check that $A_{0}$ and $A_{1}$ are as desired. ∎

### v3

@@ 3.5. Hausdorff operations and universal sets
Assume that $\mathbf{\Pi}(Z)$ has a $2^{\omega}$-universal set. The fact that $\mathbf{\Pi}(Z)$ is non-selfdual follows from Lemmas 3.35 and 3.36. To see that $\mathbf{\Pi}(Z)$ is a Wadge class, proceed as in the proof of [CMM1, Theorem 7.5]. ∎
3.6. New Wadge classes from old
Many examples of Wadge classes are obtained by combining elements of known classes using a fixed set-theoretic operation (consider for example the difference hierarchy). In fact, combining sets that belong to a given sequence of non-selfdual Wadge classes according to a given Hausdorff operation always yields a non-selfdual Wadge class (this follows from Corollary 3.39 by ignoring the sets $U_{2k+1}$). However, the operation of separated unions (to be introduced in Definition 4.5) will require in addition that certain sets are pairwise disjoint. This is why, in this subsection, the statements of our results look more complicated than one might expect.
We begin by showing that combining relativizable pointclasses according to a fixed Hausdorff operation preserves the property of having a universal set. Notice that Theorem 3.38 does not require any determinacy assumptions. The desired result on combining Wadge class will then follow easily, thanks to Theorems 3.34 and 3.37.
@@ 3.6. New Wadge classes from old
Let $D\subseteq\mathcal{P}(\omega)$, let $\mathbf{\Pi}_{n}$ for $n\in\omega$ be relativizable topological pointclasses, and let $1\leq\xi<\omega_{1}$. Assume that each $\mathbf{\Pi}_{n}$ has a $2^{\omega}$-universal set. Set
[…]
@@
The fact that $\mathbf{\Pi}$ is a relativizable pointclass is straightforward to verify using the assumption that each $\mathbf{\Pi}_{n}$ is a relativizable pointclass and Theorem 2.3. Pick a zero-dimensional space $Z$. Fix a homeomorphism $h:2^{\omega}\longrightarrow(2^{\omega})^{\omega}$. Set $(y)_{n}=\pi_{n}(h(y))$ for $y\in 2^{\omega}$ and $n\in\omega$, where $\pi_{n}:(2^{\omega})^{\omega}\longrightarrow 2^{\omega}$ denotes the projection on the $n$-th coordinate. By assumption, for every $k\in\omega$ we can fix a $2^{\omega}$-universal set $S_{2k}$ for $\mathbf{\Pi}_{k}(Z)$. Define
@@
By [Ke, Theorem 22.3], we can fix a $2^{\omega}$-universal set $U$ for $\mathbf{\Sigma}^{0}_{\xi}(Z)$. Define
@@
for $k\in\omega$, and observe that each $U_{2k+1}\in\mathbf{\Sigma}^{0}_{\xi}(2^{\omega}\times Z)$. By Theorem 2.3, there exist $U^{\prime}_{2k+1}\in\mathbf{\Sigma}^{0}_{\xi}(2^{\omega}\times Z)$ for $k\in\omega$ such that the following conditions are satisfied:
@@
will yield precisely the sequences consisting of pairwise disjoint elements of $\mathbf{\Sigma}^{0}_{\xi}(Z)$. Furthermore, changing the values of $(y)_{n}$ for even $n$ will not affect the above sequence.
@@
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $D\subseteq\mathcal{P}(\omega)$, let $Z$ be an uncountable zero-dimensional Polish space, and let $1\leq\xi<\omega_{1}$. If $\mathbf{\Gamma}_{n}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$ for $n\in\omega$ and
[…]
@@
Pick $\mathbf{\Gamma}_{n}(Z)\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$ for $n\in\omega$, where each $\mathbf{\Gamma}_{n}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$. Define relativizable pointclasses $\mathbf{\Pi}_{n}$ for $n\in\omega$ by setting $\mathbf{\Pi}_{n}(W)=\mathbf{\Gamma}_{n}(W)$ for every space $W$, and observe that each $\mathbf{\Pi}_{n}(Z)$ has a $2^{\omega}$-universal set by Theorem 3.34. It follows that $\mathbf{\Pi}(Z)$ has a $2^{\omega}$-universal set, where $\mathbf{\Pi}$ is defined as in the statement of Theorem 3.38. Since $\mathbf{\Gamma}=\mathbf{\Pi}(Z)$, where $\mathbf{\Gamma}$ is defined as in the statement of this corollary, the desired conclusion follows from Theorem 3.37. ∎
@@ 3.7. Clarifying level and expansions
In this subsection, we will show how to view level and expansions through the lens of Hausdorff operations. While the only result from this section that will be needed in the rest of the article is Proposition 3.42, the alternative descriptions given here are helpful in clarifying these two important concepts.
@@
As a first application, we will give some concrete examples of expansions. This will be useful in §5.2.
@@
We conclude this subsection with a characterization of the level that is perhaps more intuitive than the official Definition 3.19. The Borel version of Theorem 3.43 originally appeared as part of [LSR1, Théorème 8].
@@
3.8. The separation property and the “pair of socks” problem
Much like the socks in a pair, a non-selfdual Wadge class $\mathbf{\Gamma}$ and its dual $\widecheck{\mathbf{\Gamma}}$ do not appear to be distinguishable in any obvious way. Luckily, work of Van Wesep and Steel provides an elegant solution to this problem, at least when the ambient space is $\omega^{\omega}$. The key notion is the (first) separation property, already mentioned in the introduction and formally defined below.121212 The second separation property will only be needed in this subsection, and it will not appear anywhere else in this article. According to [VW1, page 92] and [VW2, page 77], both separation properties are due to Lusin. More precisely, they showed that exactly one of $\mathbf{\Gamma}$ and $\widecheck{\mathbf{\Gamma}}$ has the separation property. We will show that their result holds for uncountable zero-dimensional Polish spaces (see Theorem 3.50).
@@ 3.8. The separation property and the “pair of socks” problem
Next, we will state the original results of Van Wesep and Steel. Theorem 3.46 first appeared as [VW1, Theorem 4.2.2] (see also [VW2, Theorem 2] or [VW3, Theorem 5.3]), while Theorem 3.47 is the main result of [St].131313 The original statements of these results simply assumed $\mathsf{AD}$, but it is straightforward to check that their proofs actually yield the more precise versions given here.
Theorem 3.46 (Van Wesep).
Theorem 3.47 (Steel).
Before obtaining the desired generalization, we will need two more auxiliary results, which are essentially due to Van Wesep. In fact, Lemmas 3.48 and 3.49 are “extracted” from the proof of [VW1, Theorem 4.2.1] (see also [VW2, Theorem 1] or [VW3, Theorem 5.2]).
@@
By [Ke, Corollary 6.5], we can fix a subspace $K$ of $Z$ such that $K\approx 2^{\omega}$. Let $h:2^{\omega}\times 2^{\omega}\longrightarrow K$ be a homeomorphism. Define $j:2^{\omega}\times 2^{\omega}\times Z\longrightarrow Z\times Z$ by setting $j(x,y,z)=(h(x,y),z)$ for $(x,y,z)\in 2^{\omega}\times 2^{\omega}\times Z$, and observe that $j$ is an embedding. By Lemma 3.10, there exist $A_{0},A_{1}\in\mathbf{\Gamma}(Z\times Z)$ such that $A_{0}\cap(K\times Z)=j[A^{\prime}_{0}]$ and $A_{1}\cap(K\times Z)=j[A^{\prime}_{1}]$. It is straightforward to check that $A_{0}$ and $A_{1}$ are as desired. ∎